ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

18. Teaching and Popularization of Mathematics

Teheran, Iran 1991 MS Classification: 51 Shahvarani-Semnani, Ahmad, Shahid Beheshti Univ. , Teheran, Iran The Place of Geometry in Mathematics Currculum in Secondary Schools of Iran Nowadays the importance of teaching geometry is more evident than in previous decades, especially because as the “Discussion Document for an ICMI study” points out in recent years, research in geometry has been greatly stimulated by new ideas from inside mathematics and from other diciplines including computer sciences. In this document it is also pointed out that among mathematicians and mathematics educators there is a widespread agreement that, due to the manifold aspects of geometry, teaching of geometry should start at an early age and continue in appropriate forms throughout the whole mathematics curriculum. In Iran, geometry has been always an important part of mathematics curriculum in the educational system of the country, since the establishment of modern schools (1852). The aim of this article is to give a reasonable discussion about the place of geometry in school programs. In this study three periods have been distinguished since 1852; the period before 1955, the period between 1955-1967 and finally the period after 1967. We are here mainly concerned with the third period because in 1967 a great reform occured in the educational system in Iran and consequently important changes took place in mathematics curriculum in secondary schools. Although we will try to provide a brief discussion about the importance of geometry in the program of religious schools during the centuries before 1800. Especially we bring the views of early Iranian mathematicians towards geometry and its importance in education.